# Maxwell’s Demon at Work: Efficient Pruning by Leveraging Saturation of Neurons

### v1

@@ 6 Conclusion
In this work, we have explored how various hyperparameter configurations like the learning rate, batch size, regularization, architecture, and optimizer choices, collectively influence activation sparsity during neural network training. Leveraging this, we introduced Demon Pruning, a dynamic pruning method that controls the proliferation of saturated neurons during training through a combination of regularization and noise injection. Extensive empirical analysis on CIFAR 10 and ImageNet demonstrated superior performance compared to strong structured pruning baselines.
The simplicity of our approach allows for versatile adaptation. For settings requiring predefined sparsity levels, we can gradually increase regularization until reaching the target ratio. Furthermore, integrating our method with existing pruning techniques is straightforward. Employing multiple pruning criteria can enhance parameter identification for dead neurons, especially in unstructured methods leveraging increased structural sparsity from higher regularization.
Our experiments with Leaky ReLU illustrate the compatibility of our approach with activation functions featuring softer saturation regions than ReLU. This extends the applicability of our method to a broad range of models, offering a practical and accessible solution. Given the current trend of expanding model sizes, widespread adoption of our methodology could yield significant computational and environmental benefits.
Impact Statement
Structured pruning methods, even without specialized sparse computation primitives (Elsen et al., 2020; Gale et al., 2020), can efficiently leverage GPU hardware (Wen et al., 2016) compared to unstructured methods, which is crucial as deep learning models grow and environmental impacts escalate (Strubell et al., 2019; Lacoste et al., 2019; Henderson et al., 2020). Developing energy-efficient methods that can be widely adopted is essential.
@@ Appendix A Dead neurons in convolutional layers
In convolutional layers, ReLU is applied element-wise to the pre-activation feature map. We consider an individual neuron (filter) dead if all elements of the feature map post-activation are 0. Formally, in this case, the definition in Section 2.
Definition: The $j$-th neuron/filter in the convolutional layer $\ell$ is inactive if it consistently outputs a feature map (post-activation) with elements summing to zero on the entire training set, i.e $\sum_{k,l}F_{jkl}^{\ell}=0$. A neuron/filter that becomes and remains inactive during training is considered as dead.
@@ Appendix B Biased Random Walk Model
To formalize the intuition from section 2.1, we follow a standard line of work (Cheng et al., 2020) taking the view of SGD in Eq. 3 as a biased random walk (Anderson, 1998), described by the Langevin process,
@@
where the zero mean variable $\hat{{\bm{\xi}}}({\bm{w}},b):=\sqrt{\eta}\,(g({\bm{w}})-\hat{g}({\bm{w}},b))$ represents the gradient noise. In the limit of small learning rate, Eq. 4 is also well approximated (Cheng et al., 2020, Theorem 2) by the stochastic differential equation (SDE),
@@
In SGD, a crucial characteristic of the gradient noise is its multiplicative nature, meaning it depends on the parameter. Systems with multiplicative noise exhibit a well-known property where regions with lower noise magnitude tend to act as attractors (Oksendal, 2010). Intuitively, the noise propels the system away from regions where it has a higher impact, leading to a higher probability of staying in regions where it has a lower impact. Mathematically, this manifests as a tendency for the invariant distribution associated with SDE in Eq. (5) to have a higher probability density in regions of lower noise magnitude.
We illustrate this phenomenon using two (drastically) simplified versions of the dynamics described by Eq.( 4) and (5):
Absorbing random walk. We consider a one-dimensional absorbing random walk with a boundary at zero described by the SDE:
@@
It models a system subject to noise in an ‘active’ region $w>0$, which gets stopped at 0 – and remains there, hence ‘dies’ – once it hits 0. It can be thought of as a simplified description of a regime where the dynamics (4) is dominated by noise, such as e.g., a neuron encoding features with very low correlation to the task.
The survival probability at time $t$ is the probability that the system is still active $t$, i.e $w_{t}>0$. It is related to the distribution of the first hitting time at 0 of a standard Brownian motion, $P(T_{0}>t)$, where $T_{0}=\inf\{t\geq 0:B_{t}=0\}$. A well-known property of Brownian motion (Karatzas & Shreve, 2014) is $\lim_{t\to\infty}P(T_{0}>t)=0$, which shows that the system Eq. (6) eventually dies with probability $1$. More generally, the following proposition specifies the dependence on learning rate and initialization:
@@
[…]
@@
This is a standard application of the reflection property of Brownian motions (e.g., Lawler, 2016). Let $\bar{w}_{t}$ be the solution of Eq. (6) with no boundary condition. For the same initial condition, $w_{t}$ and $\bar{w}_{t}$ have the same distribution as long as $w_{t}>0$. Let $T_{0}=\inf\{t\geq 0:\bar{w}_{t}=0\}$ be the first hitting time of $\bar{w}_{t}$ at 00. The survival probability of $w_{t}$ can be expressed in terms of the distribution function of $T_{a}$ as
@@
[…]
Geometrical random walk. The second example illustrates the stabilizing effects of multiplicative noise for systems such as Eq. (4) near unstable critical points (Oksendal, 2010). Consider the SGD dynamics (3) with a diagonal quadratic approximation of the loss around 00, i.e., we assume
@@
where $h$ is one of the Hessian eigenvalues and $\zeta_{t}$ is a noise variable sampled from some zero mean distribution. The case of a negative eigenvalue ($h<0$) is particularly interesting since it corresponds to an unstable direction (negative curvature) in the absence of noise. In what follows, to ensure the stability of the dynamics, we assume that the noise variable $\zeta$ is bounded and the learning rate is small enough to ensure that $\eta|\zeta|<1$. We also assume the noise distribution is symmetric about 00.
@@
Combining with (14) proves Lemma B.2. ∎
@@
This is a consequence of the inequality $\log(1+x)\leq x-\frac{x^{2}}{2}+\frac{x^{3}}{3}$ for all $x>-1$. Applying this inequality to $x=-\eta(h+\zeta)$ and taking the average over $\zeta$ gives the upper bound
@@
where $\sigma^{2}:=E[\zeta^{2}]$ and we used $E[\zeta]=E[\zeta^{3}]=0$ by symmety about 00. Now, the sign of this bound coincides with the sign of the degree 2 polynomials $P(\eta):=|h|-\frac{\eta}{2}(h^{2}+\sigma^{2})+\frac{\eta^{2}}{3}|h|(h^{2}+3\sigma^{2})$. We note that $P(0)>0$ and that for a small enough ratio $|h|/\sigma$, it has two positive roots bounding an interval on which $P(\eta)<0$. One way to see this is to compute the minimum
@@
and to observe that it goes to $\infty$- ∞ as $|h|\to 0^{+}$ for fixed $\sigma^{2}$. ∎
@@ Appendix C Few Dead Neurons Revive
While empirical observations have shown a gradual accumulation of dead neurons (Fig. 1), we also observed that neurons can revive (Appendix D.2). To better assess the potential impact of reviving neurons on performance, we measured the overlap ratio ($|X\cap Y|/\min(|X|,|Y|)$) between the historical set of dead neurons at previous iterations and the set of dead neurons at the current iteration. This methodology directly follows Sokar et al. (2023). The results in figure 9 show that most neurons (over 90%) inactive at any point during training end up dead at the final iteration. This – coupled with our results showing that dying neurons can be dynamically pruned during training without impacting performance (Appendix F.2) – strongly suggests that neurons becoming inactive at any point during training in ReLU networks do not contribute significantly to the final performance of the trained model.
@@
Figure 10: Evolution of the number of dead neurons for a 3-layer MLP (with layers of widths 100, 300, and 10 respectively) on a subset of MNIST. Left: The noisy part of the minibatch gradient is isolated and used exclusively to update the parameters. Noisy updates are sufficient to kill a subset of neurons following standard initialization. Because SGD gradient is 0 for dead neurons, there is an asymmetry: only live neurons are subject to noisy updates. Center: Gaussian noise is added to the parameters update, either asymmetrically (applied only to live neurons, plain lines) or symmetrically (dashed lines). Asymmetric noise and is much more prone to dead neuron accumulation while symmetric Gaussian noise can revive neurons, contrary to SGD noise, leading to a much smaller accumulation. Right Standard SGD. Dead neurons accumulate quickly in noisy settings, but they plateau when the NN converges (leading to zero gradient). Results are averaged over 3 seeds.
@@ D.1 Noise, Learning Rate and Batch Size
To investigate the role of noise, we trained a 3 layers-deep MLP (with layers of widths 100, 300, and 10 respectively) on a subset of 10 000 images of the MNIST dataset. To isolate the noise from a minibatch (of size 1) gradient ($\hat{g}({\bm{w}}_{j}^{t})$) we subtract from it the full gradient ($g({\bm{w}}_{j}^{t})$), taken over the entire training dataset. As such, the update of neuron $j$ at every time step is given by ${\bm{w}}_{j}^{t+1}={\bm{w}}_{j}^{t}-\eta\big{(}\hat{g}({\bm{w}}_{j}^{t})-g({\bm{w}}_{j}^{t})\big{)}$.
This approach has the benefit of preserving the asymmetric noise structure of SGD updates (Wojtowytsch, 2023; Pillaud-Vivien, 2022), where dead units are not affected by noise but live ones are. Compared to applying solely symmetric Gaussian noise at every time step, we notice a much sharper accumulation of dead neurons. The details are shown in Fig. 10
Diminishing the batch size or augmenting the learning rate also creates a noisier environment because both hyperparameters affect the noise covariance in SGD optimization (Keskar et al., 2017; Masters & Luschi, 2018; Goyal et al., 2017; He et al., 2019; Li et al., 2019). Furthermore, Smith et al. (2018) shows that learning rate decay can be replaced with batch size growth, emphasizing the relationship between the two quantities. Because of their impact on noise, we should expect those quantities to affect the dying ratio of neurons. We confirm empirically this hypothesis in Fig. 11.
@@ D.3 Network Width
The widths of a neural network’s layers also influence the ratio of live neurons (live neurons to total neurons in the network) post-training (see figure 11). Typically, this ratio increases with the width; however, the total number of live neurons continues to rise with increased width. This phenomenon is somewhat anticipated as incorporating more neurons with random initialization in any given layer can only amplify the training noise, especially in the initial phase. Moreover, since initialization functions usually adjust their standard deviation proportionally to the number of channels ($\sigma\propto\sqrt{\frac{1}{\text{fan\_in}+\text{fan\_out}}}$), widening the network places neurons closer to their inactive region right from the initialization. The connection between width and dead neurons maintains its significance as neural network sizes are inclined to increase over time with the availability of more computational resources. If this trend persists, the accumulation of dead neurons could potentially become increasingly pervasive.
Figure 11: Left: An increased width leads to a higher ratio of neurons dying, independently of the optimizer. Measured on a ResNet-18 trained on CIFAR-10, without added regularization, across 3 seeds. We use the number of channels in the initial layer of the ResNet-18 to indicate the width, with 64 being the number of channels proposed by He et al. (2016) for the first convolution layer. Right: Varying the hyperparameters of a ResNet-18 (CIFAR-10) impacts the number of dead neurons. The bar heights indicate the multiplicative ratio of dead neurons compared to the base configuration ($\mathrm{lr}=0.05$, $\mathrm{bs}=128$ and $\lambda=0$). The number of training steps was kept constant when varying the batch size for fair comparisons. Quantities are averaged over 3 random seeds.
@@ Appendix E Adam is a Neuron Killer
The greater impact of Adam over the dying ratio compared to momentum must be due to the second-moment term, which is the only significant difference with momentum. Recall that Adam update is (Kingma & Ba, 2015):
@@
Earlier, we hypothesized that the neurons ending up dead were the ones experiencing very small gradients, such that the noise dominated their update trajectories. If this is the case, $g_{t}^{2}$ (the squared gradient) would be very small for those neurons’ parameters, eventually leading to a very small second-moment estimation $\hat{v}_{t}$. In such a scenario, $\epsilon$ would end up dominating $\sqrt{\hat{v}_{t}}$, effectively multiplying the learning rate by $\epsilon$ which is typically set to $1\times 10^{-8}$. Moreover, as the decay ($\beta_{2}=0.99$) of $\hat{v}_{t}$ is usually slower than the one of $\hat{m}_{t}$ ($\beta_{1}=0.9$), a few sudden noisier updates would be sufficient to make huge random steps.
It is worth noting that RL practitioners typically set epsilon to a higher value (Hessel et al., 2018), as it has empirically been found to perform better. Higher $\epsilon$ values should reduce the number of dead neurons induced by Adam optimizer, which could be the cause for the improved performance/stability observed in RL. Also, because of constant distribution shifts, rapid accumulation of dead neurons often occurs in RL tasks.
Also notable, HuggingFace Transformers library (Wolf et al., 2020) default $\epsilon$ Adam parameter to $1\times 10^{-6}$, following RoBERTa example (Liu et al., 2019). Manipulating the $\epsilon$ parameter of AdaGrad was also observed to impact significantly a transformer performance model (Agarwal et al., 2020). Verifying if those heuristic choices are due to their impact on dead neuron accumulation would be quite interesting.

### v2

@@ 5 Conclusion
In this work, we have explored how various hyperparameter configurations such as the learning rate, batch size, regularization, architecture, and optimizer choices, collectively influence activation sparsity during neural network training. Leveraging this, we introduced Demon Pruning, a dynamic pruning method that controls the proliferation of saturated neurons during training through a combination of regularization and noise injection. Extensive empirical analysis on CIFAR-10 and ImageNet demonstrated superior accuracy-sparsity tradeoffs compared to strong structured dense-to-sparse training baselines.
The simplicity of our approach allows for versatile adaptation. Integrating our method with existing pruning techniques is straightforward. Specifically, since DemP acts on the training dynamics, it can be used in conjunction with other dynamical pruning methods (evci2020rigging; Lasby2023srigl; ChaseLFSHWMMPL23). We discuss limitations of our approach in Appendix H.
Broader Impacts Statement.
Structured pruning methods, even without specialized sparse computation primitives (elsen2020fast; gale2020sparse), can efficiently leverage GPU hardware (wen2016learning) compared to unstructured methods, which is crucial as deep learning models grow and environmental impacts escalate (strubell2019energy; lacoste2019quantifying; henderson2020towards). Developing energy-efficient methods that can be widely adopted is essential.
@@
Acknowledgments
This research was enabled in part by compute resources provided by Mila, the Digital Research Alliance of Canada, and NVIDIA.
@@ Appendix A Dead neurons in convolutional layers
In convolutional layers, ReLU is applied element-wise to the pre-activation feature map. We consider an individual neuron (channel) dead if all elements in the feature map after activation are 0. Formally, in this case, the definition from Section 2 becomes:
Definition: The $j$-th neuron/channel in the convolutional layer $\ell$ is inactive if it consistently outputs a feature map (post-activation) with elements summing to zero on the entire training set, i.e $\sum_{k,l}F_{jkl}^{\ell}=0$. A neuron/channel that becomes and remains inactive during training is considered as dead.
@@ Appendix B Biased Random Walk Model
To formalize the intuition from section 2.1, we follow a standard line of work (Cheng2020_langevin) taking the view of SGD in Eq. 2 as a biased random walk (ANDERSON1998221), described by the Langevin process,
@@
where the zero mean variable $\hat{{\bm{\xi}}}({\bm{w}},b):=\sqrt{\eta}\,(g({\bm{w}})-\hat{g}({\bm{w}},b))$ represents the gradient noise. In the limit of small learning rate, Eq. 3 is also well approximated (Cheng2020_langevin, Theorem 2) by the stochastic differential equation (SDE),
@@
In SGD, a crucial characteristic of the gradient noise is its multiplicative nature, meaning it depends on the parameter. Systems with multiplicative noise exhibit a well-known property where regions with lower noise magnitude tend to act as attractors (OksendalSDE_book). Intuitively, the noise propels the system away from regions where it has a higher impact, leading to a higher probability of staying in regions where it has a lower impact. Mathematically, this manifests as a tendency for the invariant distribution associated with SDE in Eq. 4 to have a higher probability density in regions of lower noise magnitude.
We illustrate this phenomenon using two (drastically) simplified versions of the dynamics described by Eq. 3 and 4:
B.1 Absorbing random walk
We consider a one-dimensional absorbing random walk with a boundary at zero described by the SDE:
@@ B.1 Absorbing random walk
It models a system subject to noise in an ‘active’ region $w>0$, which gets stopped at 0—and remains there, hence ‘dies’—once it hits 0. It can be thought of as a simplified description of a regime where the dynamics (Eq. 3) is dominated by noise, such as e.g., a neuron encoding features with very low correlation to the task.
The survival probability at time $t$ is the probability that the system is still active $t$, i.e $w_{t}>0$. It is related to the distribution of the first hitting time at 0 of a standard Brownian motion, $P(T_{0}>t)$, where $T_{0}=\inf\{t\geq 0:B_{t}=0\}$. A well-known property of Brownian motion (KaratzasBrownianMotion_book) is $\lim_{t\to\infty}P(T_{0}>t)=0$, which shows that the system in Eq. 5 eventually dies with probability $1$. More generally, the following proposition specifies the dependence on learning rate and initialization:
@@
$P(w_{t}>0\,|\,w_{0})=\mathrm{erf}\left({\frac{w_{0}}{\sqrt{2\eta t}}}\right):=\sqrt{\frac{2}{\pi}}\int_{0}^{w_{0}/\sqrt{\eta t}}e^{\frac{-w^{2}}{2}}dw$ (6)
@@
This is a standard application of the reflection property of Brownian motions (e.g., lawler2016introduction). Let $\bar{w}_{t}$ be the solution of Eq. 5 with no boundary condition. For the same initial condition, $w_{t}$ and $\bar{w}_{t}$ have the same distribution as long as $w_{t}>0$. Let $T_{0}=\inf\{t\geq 0:\bar{w}_{t}=0\}$ be the first hitting time of $\bar{w}_{t}$ at 0. The survival probability of $w_{t}$ can be expressed in terms of the distribution function of $T_{a}$ as
@@
$P(w_{t}>0\,|\,w_{0})=1-2P(\bar{w}_{t}\leq 0)=\sqrt{\frac{2}{\pi}}\int_{0}^{w_{0}/\sqrt{\eta t}}e^{-\frac{w^{2}}{2}}dw$
B.2 Geometrical random walk
The second example illustrates the stabilizing effects of multiplicative noise for systems such as Eq. 3 near unstable critical points (OksendalSDE_book). Consider the SGD dynamics (Eq. 2) with a diagonal quadratic approximation of the loss around 0, i.e., we assume
@@ B.2 Geometrical random walk
where $h$ is one of the Hessian eigenvalues and $\zeta_{t}$ is a noise variable sampled from some zero mean distribution. The case of a negative eigenvalue ($h<0$) is particularly interesting since it corresponds to an unstable direction (negative curvature) in the absence of noise. In what follows, to ensure the stability of the dynamics, we assume that the noise variable $\zeta$ is bounded and the learning rate is small enough to ensure that $\eta|\zeta|<1$. We also assume the noise distribution is symmetric about 0.
@@
Combining with Eq. 13 proves Lemma B.2. ∎
@@
This is a consequence of the inequality $\log(1+x)\leq x-\frac{x^{2}}{2}+\frac{x^{3}}{3}$ for all x>1x>-1. Applying this inequality to $x=-\eta(h+\zeta)$ and taking the average over $\zeta$ gives the upper bound
@@
where $\sigma^{2}:=E[\zeta^{2}]$ and we used $E[\zeta]=E[\zeta^{3}]=0$ by symmety about 0. Now, the sign of this bound coincides with the sign of the degree 2 polynomials $P(\eta):=|h|-\frac{\eta}{2}(h^{2}+\sigma^{2})+\frac{\eta^{2}}{3}|h|(h^{2}+3\sigma^{2})$. We note that $P(0)>0$ and that for a small enough ratio $|h|/\sigma$, it has two positive roots bounding an interval on which $P(\eta)<0$. One way to see this is to compute the minimum
@@
and to observe that it goes to -$\infty$ as $|h|\to 0^{+}$ for fixed $\sigma^{2}$. ∎
@@ Appendix C Few Dead Neurons Revive
While empirical observations have shown a gradual accumulation of dead neurons (Fig. 1), we also observed that neurons can revive (Appendix D.2). To better assess the potential impact of reviving neurons on performance, we measured the overlap ratio ($|X\cap Y|/\min(|X|,|Y|)$) between the historical set of dead neurons at previous iterations and the set of dead neurons at the current iteration. This methodology directly follows sokar2023dormant. The results in Fig. 8 show that most neurons (over 90%) inactive at any point during training end up dead at the final iteration. This—coupled with our results showing that dying neurons can be dynamically pruned during training without impacting performance (Appendix F.2)—strongly suggests that neurons becoming inactive at any point during training in ReLU networks do not contribute significantly to the final performance of the trained model.
@@
Figure 9: Evolution of the number of dead neurons for a 3-layer MLP (with layers of widths 100, 300, and 10 respectively) on a subset of MNIST. Left: The noisy part of the minibatch gradient is isolated and used exclusively to update the parameters. Noisy updates are sufficient to kill a subset of neurons following standard initialization. Because SGD gradient is 0 for dead neurons, there is an asymmetry: only live neurons are subject to noisy updates. Center: Gaussian noise is added to the parameters update, either asymmetrically (applied only to live neurons, plain lines) or symmetrically (dashed lines). Asymmetric noise is much more prone to dead neuron accumulation while symmetric Gaussian noise can revive neurons, contrary to SGD noise, leading to a much smaller accumulation. Right Standard SGD. Dead neurons accumulate quickly in noisy settings, but they plateau when the NN converges (leading to zero gradient). Results are averaged over three seeds.
@@ D.1 Noise, Learning Rate and Batch Size
To investigate the role of noise, we trained a 3-layer deep MLP (with layers of widths 100, 300, and 10 respectively) on a subset of 10 000 images of the MNIST dataset. To isolate the noise from a minibatch (of size 1) gradient ($\hat{g}({\bm{w}}_{j}^{t})$) we subtract from it the full gradient ($g({\bm{w}}_{j}^{t})$), taken over the entire training dataset. As such, the update of neuron $j$ at every time step is given by ${\bm{w}}_{j}^{t+1}={\bm{w}}_{j}^{t}-\eta\big(\hat{g}({\bm{w}}_{j}^{t})-g({\bm{w}}_{j}^{t})\big)$.
This approach has the benefit of preserving the asymmetric noise structure of SGD updates (DBLP:journals/jns/Wojtowytsch23; Pillaud-Vivien_blog), where dead units are not affected by noise but live ones are. Compared to applying solely symmetric Gaussian noise at every time step, we notice a much sharper accumulation of dead neurons. The details are shown in Fig. 9.
Diminishing the batch size or augmenting the learning rate also creates a noisier environment because both hyperparameters affect the noise covariance in SGD optimization (DBLP:conf/iclr/KeskarMNST17; DBLP:journals/corr/MastersLuschi2018; DBLP:journals/corr/GoyalDGNWKTJH17; DBLP:conf/nips/HeLT19; DBLP:conf/nips/LiWM19a). Furthermore, DBLP:conf/iclr/SmithKYL18 shows that learning rate decay can be replaced with batch size growth, emphasizing the relationship between the two quantities. Because of their impact on noise, we should expect those quantities to affect the dying ratio of neurons. We confirm empirically this hypothesis in Fig. 10.
@@ D.2 Training Time
Figure 10: Left: An increased width leads to a higher ratio of neurons dying, independently of the optimizer. Measured on a ResNet-18 trained on CIFAR-10, without added regularization, across three seeds. We use the number of channels in the initial layer of the ResNet-18 to indicate the width, with 64 being the number of channels proposed by DBLP:conf/cvpr/HeZRS16 for the first convolution layer. Right: Varying the hyperparameters of a ResNet-18 (CIFAR-10) impacts the number of dead neurons. The bar heights indicate the multiplicative ratio of dead neurons compared to the base configuration ($\mathrm{lr}=0.05$, $\mathrm{bs}=128$ and $\lambda=0$). The number of training steps was kept constant when varying the batch size for fair comparisons. Quantities are averaged over three random seeds.
@@ D.3 Network Width
The widths of a neural network’s layers also influence the ratio of live neurons (live neurons to total neurons in the network) post-training (see Fig. 10). Typically, this ratio increases with the width; however, the total number of live neurons continues to rise with increased width. This phenomenon is somewhat anticipated as incorporating more neurons with random initialization in any given layer can only amplify the training noise, especially in the initial phase. Moreover, since initialization functions usually adjust their standard deviation proportionally to the number of channels ($\sigma\propto\sqrt{\frac{1}{\text{fan\_in}+\text{fan\_out}}}$), widening the network places neurons closer to their inactive region right from the initialization. The connection between width and dead neurons maintains its significance as neural network sizes are inclined to increase over time with the availability of more computational resources. If this trend persists, the accumulation of dead neurons could potentially become increasingly pervasive.
@@ Appendix E Adam is a Neuron Killer
The greater impact of Adam over the dying ratio compared to momentum must be due to the second-moment term, which is the only significant difference with momentum. Recall that Adam update (DBLP:journals/corr/KingmaB14) is given by:
@@
Earlier, we hypothesized that the neurons ending up dead were the ones experiencing very small gradients, such that the noise dominated their update trajectories. If this is the case, $g_{t}^{2}$ (the squared gradient) would be very small for those neurons’ parameters, eventually leading to a very small second-moment estimation $\hat{v}_{t}$. In such a scenario, $\epsilon$ would end up dominating $\sqrt{\hat{v}_{t}}$, effectively multiplying the learning rate by $\epsilon^{-1}$ ($\epsilon$ is typically set to $1\times 10^{-8}$). Moreover, as the decay ($\beta_{2}=0.99$) of $\hat{v}_{t}$ is usually slower than the one of $\hat{m}_{t}$ ($\beta_{1}=0.9$), a few sudden noisier updates would be sufficient to make huge random steps.
It is worth noting that RL practitioners typically set epsilon to a higher value (DBLP:conf/aaai/HesselMHSODHPAS18), as it has empirically been found to perform better. Because of constant distribution shifts, rapid accumulation of dead neurons often occurs in RL tasks (lyle2022understanding; sokar2023dormant). Higher $\epsilon$ values should reduce the number of dead neurons induced by Adam optimizer, by making the effective learning rate ($\frac{\eta}{\epsilon}$) smaller when $\epsilon$ dominates $\sqrt{\hat{v}_{t}}$. We confirm this hypothesis in Fig. 11, confirming the soundness of picking higher $\epsilon$ to improve performance and stability in RL.
Figure 11: Sparsity decreases from 5.07±2.04% to 0% when varying Adam $\epsilon$ parameter from 1e-10 to 1e-4, or from 12.6%±1.15% to 0.47%±0.01 when the learning rate is increased, supporting our claim that the usage Adam optimizer impacts the accumulation of dead neurons through $\epsilon$ parameter. We note that this additional evidence complements the findings from lyle2022understanding, which reported reduced plasticity loss with increasing $\epsilon$. Experiments were performed with ResNet-18 on CIFAR-10 over three seeds, without DemP, using the same training recipe as in Appendix I.1.
Also notable, HuggingFace Transformers library (wolf-etal-2020-transformers) default $\epsilon$ Adam parameter to $1\times 10^{-6}$, following RoBERTa example (DBLP:journals/corr/Roberta2019). Manipulating the $\epsilon$ parameter of AdaGrad was also observed to impact significantly a transformer performance model (DBLP:journals/corr/AgarwalAHKZ2020). Confirming that those heuristic choices are due to their impact on dead neuron accumulation would be quite interesting.